OpenStudy (ayee_ciera):

which represents the solution to the absolute value equation? 3|2x+4|-1=11

OpenStudy (lynfran):

3|2x+4|-1=11 3|2x+4|=11+1 3|2x+4|=12 |2x+4|=12/3 |2x+4|=4 (2x+4)=+4 and (2x+4)=-4 can find solution frm here...?

OpenStudy (ayee_ciera):

yes thank you @LynFran

OpenStudy (lynfran):

welcome

OpenStudy (ayee_ciera):

@LynFran the answers i got doesnt match the answer choices

OpenStudy (lynfran):

2x+4=4 2x=4-4 2x=0 x=0 and 2x+4=-4 2x=-4-4 2x=-8 x=-8/2 x=-4

OpenStudy (ayee_ciera):

ohhh i see where i messed up thank you @LynFran

OpenStudy (lynfran):

welcome